How large a force parallel to a 300 incline is needed to give a 5.0 kg an acceleration 0.20 m/s2 up the incline a) if friction is negligible? b) if the coefficient is 0.30?


1
Expert's answer
2022-04-15T16:08:04-0400

a)

"F=m(a+g\sin30°)=26~N."

b)

"F=m(a+g\sin30°+\mu g\cos30°)=39~N."